UNLV CS456 Regular Languages

Name:

1. Give a regular expression for each of these languages. (a) All binary strings which do not contain ‘000’ as a substring.

(b) Base 5 numerals for positive multiples of 3, where leading zeros are not allowed.

2. Find a minimal DFA equivalent to each of these NFAs. (a)

(b) 3 (c)

4 (d) 3.

5